OpenStudy (anonymous):

a rectangular table is two times as long as it is wide. If the area is 72 ft ^2, find the length and width of the table

Directrix (directrix):

Let w = width Then, the length L is 2w Area of a rectangle = L*W A = (2w)*w 72 = 2 w ^2

Directrix (directrix):

Will you do the next step?

OpenStudy (anonymous):

to be honest, i'm not sure how but I can try

OpenStudy (anonymous):

72=2w^2 72/2=2w^2/2 36=w^2?

Directrix (directrix):

Right. Take square roots of both sides. Remember that in an equation such as x^2 = 4, x = 2 or -2.

OpenStudy (anonymous):

okay so then w=18?

Directrix (directrix):

Taking half is not the same as square rooting.

OpenStudy (anonymous):

oh its 6 isn't it

OpenStudy (anonymous):

okay well my name says it all...

Directrix (directrix):

@Isuckatmath1 -->6 and -6 but -6 has no meaning in this context.

Directrix (directrix):

You are correct in the answer but wrong at sucking in math.

OpenStudy (anonymous):

the answer says it's wrong

OpenStudy (anonymous):

Area A = L * w ->72 = 2w * w = 2 w^2 -> w^2 = 72/2 = 36 => w = sqrt ( 36) = 6 ft

Directrix (directrix):

Did you include the length = 2 w?

Directrix (directrix):

W = 6 and then L = the length and width of the table

OpenStudy (anonymous):

yeah i don't know what I am doing wrong but I have to just post it because I only have 3 mins left to finish. Thank you both for trying to help me though.

OpenStudy (anonymous):

Did you post

OpenStudy (anonymous):

it says that w = 6 and l = 12

OpenStudy (anonymous):

L = 2 w = 12 ft?

Directrix (directrix):

that's what we got.

OpenStudy (anonymous):

I thought we got 36 and 6

OpenStudy (anonymous):

@Directrix, Medal for being consistently patient :P

OpenStudy (anonymous):

Isuck, READ, READ, READ................. at least the last sentence! w (width) = 6

OpenStudy (anonymous):

I did read, I am bad at math not stupid...

OpenStudy (anonymous):

I don't say this word, I feel you're way too panic, so you overlook the last result!